Status Report

NOAA GOES SATOPS Morning Report 10 Jan 2001

By SpaceRef Editor
January 10, 2001
Filed under ,

SATOPS MORNING REPORT – January 10, 2001

Human error caused a total failure of the tech bus to the Wallops CDA
Operations Building during annual maintenance of the switchgear equipment. 
Recovery took from approximately 1430Z to 1540Z for GOES imaging. 
EMWIN,WEFAX, were restored by 1700Z, METEOSAT-7 was restored at 1730Z. 
Wallops was able to capture the POES (NOAA-11) data in a stand-alone mode
(13M antenna and IMUX powered, PACS powered but not up) without any impact
to operations.

We were quite fortunate to only have minor equipment problems (METEOSAT
modem, EMWIN, WEFAX ) from a hard crash.  Switchgear maintenance is
expected to resume today.

POES

NOAA-16 experienced an anomalous two count jump in the AVHRR sync delta. 
This “hiccup” is under analysis and monitoring by engineering and operations.

NOAA-14 rev 31976F 1807Z lost 14% of expected data due to signal fade
during the playback.  A personal error in configuring for a repeat
playback on rev 31078W 2112Z resulted in a more marginal playback than
the original, recorder management precluded a third attempt.

NOAA-16 rev 1560F 0056Z System “A” mid S-band combiner problems were
overcome by crew replaying G1A a second time during the contact. 
G3A was noisy (82% recovered due to dropouts) and was rescheduled and recovered
completely on rev 1564W 0805Z.

NOAA-14 rev 31081F 0225Z System “A” low S-band combiner problems were
overcome by crew replaying G5A a second time during the contact.

NOAA-15 rev 13831F 0314Z System “A” emergency S-band combiner problems
were overcome by crew replaying G1A a second time during the contact.

Fairbanks technicians are investigating the System “A” combiner problems.

 

DMSP

OPS48 rev 19439F System “B” failed to go into pass mode, System “C”
was brought on-line in time to accomplish processor tasks but not in time
to recover data.  Crew added an additional contact to accomplish data
recovery.

OPS49 violated the red limit threshold on 2VEAX(Y)(Z), engineering support
was notified at 0600Z.  Point are nominal this morning.

 

GOES

No change in the status of the GOES spacecraft.

GOES East:  All data lost at 1425Z due to tech bus power interruption
at Wallops CDA.  Sounder resumed operations at 1520Z, imager resumed
operations at 1539Z without Image Motion Conpensation (IMC).  IMC
enabled for 1601Z image.  Addition IMC and boomsnap problems at 1637Z
were cleared at the expense of the 1640Z image.  EMWIN/WEFAX restored
at 1700Z.

GOES West:  All data lost at 1425Z due to tech bus power interruption
at Wallops CDA.  Recovered from a “stuck in star” situation and resumed
post maneuver recovery operations with the 1531Z image.  Ground system
orbit and attitude determination globals restored for the 1601Z image. 
EMWIN/WEFAX restored at 1700Z.

 

Other: Total lunar eclipse at 1950Z today (2:50pm local) and Wallops
power system breaker testing occuring this morning.

 

Questions or comments to [email protected]

SpaceRef staff editor.